# Bielik v3 Small: Technical Report

### v1

@@ Abstract
We introduce Bielik v3, a series of parameter-efficient generative text models (1.5B and 4.5B) optimized for Polish language processing. These models demonstrate that smaller, well-optimized architectures can achieve performance comparable to much larger counterparts while requiring substantially fewer computational resources. Our approach incorporates several key innovations: a custom Polish tokenizer (APT4) that significantly improves token efficiency, Weighted Instruction Cross-Entropy Loss to balance learning across instruction types, and Adaptive Learning Rate that dynamically adjusts based on training progress. Trained on a meticulously curated corpus of 292 billion tokens spanning 303 million documents, these models excel across multiple benchmarks, including the Open PL LLM Leaderboard, Complex Polish Text Understanding Benchmark, Polish EQ-Bench, and Polish Medical Leaderboard. The 4.5B parameter model achieves results competitive with models 2-3 times its size, while the 1.5B model delivers strong performance despite its extremely compact profile. These advances establish new benchmarks for parameter-efficient language modeling in less-represented languages, making high-quality Polish language AI more accessible for resource-constrained applications.
@@ 1 Introduction
Several notable efforts have advanced Polish language modeling in recent years. TRURL 2, a collection of fine-tuned Llama 2 models with 7 billion and 13 billion parameters, was trained on approximately 1 million conversational samples. Qra models, comprising continuously pretrained architectures with 1, 7, and 13 billion parameters, leveraged 90 billion tokens of Polish data. More recently, PLLuM, developed by a consortium of Polish academic institutions, introduced models ranging from 8 billion to 70 billion parameters, created through continued pretraining of Llama and Mistral models on Polish corpora. While these initiatives have made important strides, they often face limitations in performance, versatility, or accessibility, frequently requiring significantly larger computational resources for comparable performance.
Building on our previous work with Bielik 7B v0.1 and Bielik 11B v2, we introduce the Bielik v3 series of generative text models optimized specifically for Polish language processing. These models, with sizes of 1.5B and 4.5B parameters, represent a significant advancement in parameter-efficient language modeling. By adopting an innovative approach to model scaling and training, we demonstrate that smaller, well-optimized models can achieve performance comparable to much larger counterparts while requiring substantially fewer computational resources.
Our approach introduces several key technical innovations. First, we employ depth up-scaling to adapt Qwen2.5 models, replacing the original tokenizer with a custom-developed Polish tokenizer (APT4) that significantly improves token efficiency for Polish texts. Second, we implement Weighted Instruction Cross-Entropy Loss, which balances the learning of different instruction types by assigning quality-based weights to training examples. Third, we utilize Adaptive Learning Rate, which dynamically adjusts the learning rate based on training progress and context length. These techniques, combined with comprehensive training on a diverse corpus of 292 billion tokens across 303 million documents, enable the Bielik v3 models to achieve remarkable performance despite their compact size.
@@ 5.5 Reinforcement Learning
Group Relative Policy Optimization (GRPO) is a reinforcement learning algorithm designed to fine-tune large language models (LLMs) by optimizing their policy based on group-relative advantages. Unlike traditional methods like Proximal Policy Optimization (PPO), which rely on a separate value function to estimate the expected return, GRPO eliminates the need for this component, thereby reducing memory consumption and computational complexity.
@@
Optimization settings featured a learning rate of 1e-6, mini-batch size of 128, and micro-batch size per GPU of 16. To ensure stable policy updates, KL divergence regularization was applied with a coefficient of 0.001 using the low-variance KL loss type.
@@ 8 Conclusion
    Training Methodology Innovations: Our techniques include Weighted Instruction Cross-Entropy Loss and Adaptive Learning Rate, which significantly improved model performance, particularly for Polish-specific linguistic patterns.

### v2

@@ Abstract
We introduce Bielik v3, a series of parameter-efficient generative text models (1.5B and 4.5B) optimized for Polish language processing. These models demonstrate that smaller, well-optimized architectures can achieve performance comparable to much larger counterparts while requiring substantially fewer computational resources. Our approach incorporates key innovations, including a custom Polish tokenizer (APT4) that significantly improves token efficiency and Adaptive Learning Rate that dynamically adjusts based on training progress. Trained on a meticulously curated corpus of 292 billion tokens spanning 303 million documents, these models excel across multiple benchmarks, including the Open PL LLM Leaderboard, Complex Polish Text Understanding Benchmark, Polish EQ-Bench, and Polish Medical Leaderboard. The 4.5B parameter model achieves results competitive with models 2-3 times its size, while the 1.5B model delivers strong performance despite its extremely compact profile. These advances establish new benchmarks for parameter-efficient language modeling in less-represented languages, making high-quality Polish language AI more accessible for resource-constrained applications.
@@ 1 Introduction
Several notable efforts have advanced Polish language modeling in recent years. TRURL 2 Voicelab (2023), a collection of fine-tuned Llama 2 models with 7 billion and 13 billion parameters, was trained on approximately 1 million conversational samples. Qra National Information Processing Institute and Gdańsk University of Technology (2024) models, comprising continuously pretrained architectures with 1, 7, and 13 billion parameters, leveraged 90 billion tokens of Polish data. More recently, PLLuM, developed by a consortium of Polish academic institutions, introduced models ranging from 8 billion to 70 billion parameters, created through continued pretraining of Llama and Mistral models on Polish corpora. While these initiatives have made important strides, they often face limitations in performance, versatility, or accessibility, frequently requiring significantly larger computational resources for comparable performance.
Building on our previous work with Bielik 7B v0.1 Ociepa et al. (2024) and Bielik 11B v2 Ociepa et al. (2025), we introduce the Bielik v3 series of generative text models optimized specifically for Polish language processing. These models, with sizes of 1.5B and 4.5B parameters, represent a significant advancement in parameter-efficient language modeling. By adopting an innovative approach to model scaling and training, we demonstrate that smaller, well-optimized models can achieve performance comparable to much larger counterparts while requiring substantially fewer computational resources.
Our approach introduces several key technical innovations. First, we employ depth up-scaling to adapt Qwen2.5 models Qwen et al. (2025), replacing the original tokenizer with a custom-developed Polish tokenizer (APT4) that significantly improves token efficiency for Polish texts. Second, we utilize Adaptive Learning Rate, which dynamically adjusts the learning rate based on training progress and context length. These techniques, combined with comprehensive training on a diverse corpus of 292 billion tokens across 303 million documents, enable the Bielik v3 models to achieve remarkable performance despite their compact size.
@@ 5.5 Reinforcement Learning
The optimization settings included a learning rate of $1\times 10^{-6}$, a global batch size of 128 (with a local batch size of 16 per GPU). To ensure stable policy updates, KL divergence regularization was applied with a coefficient of 0.001, using the low-variance KL loss type.
@@ 6.4 Polish Medical Leaderboard
6.5 Polish Linguistic and Cultural Competency Benchmark (PLCC)
The Polish Linguistic and Cultural Competency Benchmark (PLCC) (Dadas et al., 2025) is a specialized evaluation framework that tests language models’ grasp of Polish cultural knowledge and context. Unlike conventional NLP assessments, PLCC delves deeper into cultural understanding through 600 carefully designed questions spanning six key domains: history, geography, culture & tradition, art & entertainment, grammar, and vocabulary.
The benchmark’s questions probe models’ familiarity with various aspects of Polish heritage, including cultural references, historical milestones, traditional customs, folklore, literary works, and contemporary popular culture. These elements are crucial for achieving authentic language comprehension that goes beyond mere text processing. The questions vary in complexity from widely known information to specialized regional knowledge, presented in both multiple-choice and open-ended formats that demand precise responses with specific facts, dates, names, or concepts.
Model Parameters (B) Average Score (%)
Gemini-2.5-Pro-Exp-03-25 Unknown 89.50
DeepSeek-R1 685.0 76.00
DeepSeek-v3-0324 685.0 71.00
DeepSeek-v3 685.0 69.17
PLLuM-8x7B-nc-chat 46.7 68.17
Llama-3.1-Tulu-3-405B 405.0 63.83
Bielik-11B-v2.2-Instruct 11.2 63.00
Bielik-11B-v2.3-Instruct 11.2 62.17
GPT-4.1-mini-2025-04-14 Unknown 62.17
Bielik-11B-v2.1-Instruct 11.2 61.00
Llama-3.1-405B 405.0 60.00
PLLuM-12B-nc-chat 12.2 59.50
Llama-PLLuM-70B-chat 70.6 58.50
Llama-4-Maverick 402.0 58.17
Command-A-03-2025 111.0 56.17
Mistral-Large-2407 123.0 54.17
PLLuM-8x7B-chat 46.7 54.17
Mistral-Large-2411 123.0 52.00
WizardLM-2-8x22B 141.0 51.50
Qwen-Max Unknown 50.83
Command-R-Plus-08-2024 Unknown 50.17
Mixtral-8x22B 141.0 49.83
Command-R-Plus-04-2024 Unknown 49.33
Llama-3.3-70B 70.6 48.83
Llama-3.1-70B 70.0 47.83
Gemma-3-27B 27.4 47.33
PLLuM-12B-chat 12.2 47.00
Bielik-7B-Instruct-v0.1 7.0 46.67
Mistral-Small-3.1-24B-2503 24.0 43.33
Llama-3.0-70B 70.0 43.00
Gemma-2-27B 27.0 42.67
Bielik-4.5B-v3.0-Instruct 4.8 42.33
Llama-4-Scout 109.0 41.50
EuroLLM-9B 9.0 41.00
Qwen-2.5-72B 72.7 39.17
Mistral-Small-24B-2501 24.0 39.00
Llama-PLLuM-8B-chat 8.0 38.50
Qwen3-32B 32.8 37.67
Mixtral-8x7B 46.7 35.33
Qwen-2.5-32B 32.8 30.50
Qwen3-14B 14.8 30.33
Gemma-2-9B 9.0 29.17
Phi-4 14.7 29.17
Bielik-1.5B-v3.0-Instruct 1.6 27.50
Qwen-2.5-14B 14.8 26.67
Qwen3-8B 8.2 26.00
Mistral-Nemo 12.2 23.00
Command-R-7B 7.0 22.83
Llama-3.1-8B 8.0 22.67
Mistral-7B-v0.3 7.2 21.83
Ministral-8B 8.0 20.67
Qwen-2.5-7B 7.0 17.67
Models with a non-commercial license.
Table 14: Polish Linguistic and Cultural Competency Benchmark (PLCC) results for open-source models. Closed proprietary models have been excluded from this comparison.
The PLCC results (Table 14) reveal several insights about Bielik v3 models’ cultural competency:
Parameter efficiency:
Bielik-4.5B-v3.0-Instruct achieves 42.33% on PLCC despite its small size, outperforming several larger models including Qwen-2.5-14B (26.67%) and Phi-4 (29.17%). Similarly, Bielik-1.5B-v3.0-Instruct reaches 27.50%, demonstrating strong performance for its compact parameter count.
Size-performance tradeoff:
While Bielik v3 models score lower than their larger Bielik v2 counterparts (Bielik-11B-v2.2-Instruct at 63.00%), they provide viable alternatives for resource-constrained deployments while maintaining reasonable cultural knowledge capabilities.
Competitive small-model performance:
Bielik-4.5B-v3.0-Instruct shows competitive results compared to models with similar parameter counts, highlighting the effectiveness of our training methodology in preserving cultural knowledge despite parameter reduction.
Progressive improvements:
The performance gap between Bielik-1.5B-v3.0-Instruct (27.50%) and Bielik-4.5B-v3.0-Instruct (42.33%) demonstrates the benefits of additional parameters for cultural knowledge retention, while both maintain efficient footprints.
@@ 8 Conclusion
    Training Methodology Innovations: Our techniques include Adaptive Learning Rate, which significantly improved model performance, particularly for Polish-specific linguistic patterns.